Two liq A and B have vapour pressure of 500 mm Hg and 200 mm Hg respect. calculate the mole fraction of A at which two liquids have equal partial pressure.

Grade:12

1 Answers

Yash Jain
9 years ago
Given that, P= PB
So, xAPoA  = xBPoB
or, nA*500 = nB*200
gives, nA/nB = 2/5
thus, nA = 2x & nB = 5x  {for any integer x}
Then, xA = nA / nA+nB = 2x / 2x+5x = 2/7
and xB = 1 - xA = 5/7